OpenStudy (anonymous):

determine the intervals over which the function is increasing y=2x^3+3x^2-12x

OpenStudy (anonymous):

Can you take the derivative??

OpenStudy (anonymous):

Take the derivative and put that is equal to 0..

OpenStudy (anonymous):

didnt learn about derivatives yet :(

OpenStudy (anonymous):

:) then, compute \[\frac{f(x)-f(y)}{x-y}\] and see where It's positive or negative !

OpenStudy (anonymous):

\[ f(x)=2 x^3+3 x^2-12 x\\ f'(x)=-12 + 6 x + 6 x^2=\\ 6 (-1 + x) (2 + x)\\ f'(x) >0 \text { when } x < -2 \text { or } x> 1\\ f'(x) <0 \text { when } -2 < x <1 \] f is increasing on \( ]-\infty -2[\cup]1, \infty[\) f is decreasing on \( ]-2, 1[\)